Mr. and Mrs. Daku have the following income items. Mr. Dakus salary Mrs. Dakus Schedule C net profit Interest income $ 64,7Married filing jointly and surviving spouses Married filing separately Head of household Single $24,400 12,200 18,350 12,200Individual Tax Rate Schedules Married Filing Jointly and Surviving Spouse If taxable income is: The tax is: Not over $19,400
Answer #1

AGI = Mr. Daku’s salary+Mrs. Daku’s Schedule C net profit+Interest income-(1/2 * Mrs. Daku’s self-employment tax) = 64750+45200+2500-(1/2*6387) = $109257

Taxable income = AGI-Standard deduction – Section 199A deduction = 109257-24400-(20%*45200) = $75817 (standard deduction for married filing jointly in 2019 is $24400)

Income tax = 1940+(12%*(75817-19400)) = $8710 (using tax rate schedule of married filing jointly)